OpenStudy (aleah54):

Simplify (3xy^2)(4xy)(2xy)^3. Can you explain how you got your answer

OpenStudy (gunsodo):

\[(3xy^2)(4xy)(2xy)^3 = (3\times4\times2^3)xy^2 \times xy\times x^3y^3=96x^5y^6\]

OpenStudy (aleah54):

where did you get the 3 when you typed x^3 y^3

OpenStudy (aleah54):

so you mutiple 3 x 4 x 2^3 you get 60 what is the next step

OpenStudy (phi):

now do the letter "x" if we focus only on the x's you have x*x*x^3 can you simplify that ?

OpenStudy (aleah54):

would it be just 3x

OpenStudy (phi):

no. 3x means x+x+x we have x*x*x^3 (remember, x^3 is "short-hand" for x*x*x) so you really have x*x*x*x*x people use "short-hand" to write that do you know how to write it using the "short" way ?

OpenStudy (aleah54):

no

OpenStudy (phi):

in x*x*x*x*x how many times is x being multiplied by itself ?

OpenStudy (phi):

in other words, how many x's do you see ?

OpenStudy (aleah54):

4

OpenStudy (phi):

x*x*x*x*x that has 5 x's so the short way to write it is x with a little 5 in the upper right \[ x^5\] or write x^5 if we don't use the equation editor that means so far your answer is 3*4*2^3 which is 3*4*2*2*2 or 3*4*8 or 12*8 or 96 and x*x*x*x*x or x^5 or 96x^5 now we do the y's

OpenStudy (phi):

the y's are y^2 * y * y^3

OpenStudy (aleah54):

?

OpenStudy (phi):

you could do it the long way what is y^2 written the long way ? It is y*y what is y^3 written the long way? do you know ?

OpenStudy (aleah54):

y y y

OpenStudy (phi):

yes, so y^2 * y * y^3 is also written the long way as y*y*y* y*y*y now write that the short-way can you do that ?

OpenStudy (aleah54):

yes it would be y y and just y y y right

OpenStudy (phi):

the "short-way" means putting a little number in the upper right of y what number should we use for y*y*y*y*y*y ?

OpenStudy (aleah54):

6

OpenStudy (phi):

yes, so instead of y*y*y*y*y*y we write \(y^6\) or y^6 (It's easier to use the short-way, right ?)

OpenStudy (aleah54):

yes

OpenStudy (phi):

so we write y^2 * y * y^3 as y^6

OpenStudy (phi):

and the simplified way to write (3xy^2)(4xy)(2xy)^3 is 96 x^5 y^6

OpenStudy (phi):

btw, the trickiest part is (2xy)^3 that is short for writing (2xy)(2xy)(2xy) or 2*x*y*2*x*y*2*x*y when you multiply, you can change the order, so you can write that mess as 2*2*2*x*x*x*y*y*y or 2^3 x^3 y^3